Queens, Inc. has a total debt ratio of .46. what is its Debt-equity ratio? What is its equity multiplier?

Please show steps for

Total debt ratio
Total assets
Total equity
total debt ratio
Debt to equity ratio
equity multiplier
Answer #1

Answers:- 1) 0.85 2)1.85

1)debt equity ratio

Debt ratio = debt/assets

(debt+equity) = assets { (always it shows 100%) }

Debt ratio = debt / (debt+equity)

0.46 = debt/(debt+equity)

0.46equity =1debt-0.46 debt

0.54debt =0.46 equity

debt/equity ratio =0.85

2) Equity multiplier = 1+ (debt/equity)

=1+0.85 = 1.85